### v2

# Asymgenesis

Abstract
We present a framework based on the standard type-I seesaw model that relates the baryon asymmetry of the universe to the dark matter (DM) density. The framework, which we name Asymgenesis, relies on the presence of primordial charge asymmetries seeded either in the dark sector or in the visible sector. A higher-dimensional portal operator reshuffles this initial asymmetry into both sectors, eventually resulting in a nonzero $B-L$ asymmetry and an asymmetric DM component. Compared to conventional asymmetric-dark-matter (ADM) schemes, our framework imposes far milder requirements on the portal interaction. In particular, the portal interaction need not violate $B-L$, and the temperature scales of efficient $B-L$ violation and efficient charge-transfer interaction mediated by the portal operator can be separated. We develop the formalism in detail and argue that the flexibility of our framework enlarges the model-building landscape for ADM.
preprint: MITP-25-051
Introduction —  Addressing the observed baryon asymmetry of the Universe (BAU) Planck:2018vyg ; ParticleDataGroup:2024cfk calls for physics beyond the Standard Model (SM). A leading paradigm for dynamically generating the BAU in the early universe is leptogenesis (LG) Fukugita:1986hr , which requires extending the SM field content by at least two right-handed neutrinos (RHNs). An attractive feature of LG is that RHNs can account for the observed nonzero masses of active neutrinos via the type-I seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1980xy ; Gell-Mann:1979vob ; Mohapatra:1979ia , thereby connecting the origin of the BAU to the origin of active neutrino masses Buchmuller:2005eh ; Bodeker:2020ghk . Most LG realizations rely on all three of Sakharov’s conditions Sakharov:1967dj for generating the BAU to emerge from the RHN sector Fukugita:1986hr ; Luty:1992un ; Covi:1996wh ; Pilaftsis:1997jf ; Akhmedov:1998qx ; Barbieri:1999ma ; Pilaftsis:2003gt ; Giudice:2003jh ; Buchmuller:2004nz ; Pilaftsis:2005rv ; Klaric:2020phc ; Klaric:2021cpi . A notable exception is wash-in leptogenesis (WILG), which does not require any source of charge-parity $(CP)$-violation from the RHN sector Domcke:2020quw ; Domcke:2022kfs ; Mojahed:2025vgf ; WILGRelated .
The philosophy underlying WILG is to treat RHN interactions as spectator processes whose role is to provide a source of $B\!-\!L$ violation. In this way, RHNs are treated on par with electroweak sphalerons Kuzmin:1985mm , whose role in standard LG mechanisms is merely to serve as a source of $B+L$ violation. Meanwhile, $CP$ violation is attributed to novel $CP$-violating dynamics at high temperatures above the mass of the heaviest RHN $T\gg M_{N}$, in a process referred to as chargegenesis (CG), see Refs. Domcke:2022kfs ; Schmitz:2023pfy ; Mukaida:2024eqi ; Mojahed:2024yus for concrete realizations. The role of CG is to provide $CP$-violating initial conditions for WILG, and can be viewed as a UV completion for the effective framework of WILG. In particular, the SM entails eleven linearly independent conserved global charges at high temperatures that could be produced during CG Domcke:2020quw . During WILG, RHNs act on the nontrivial chemical background induced by CG in a $B-L$ violating manner and drive the thermal plasma to an attractor solution featuring nonzero $B-L$, even if $B-L$ was conserved throughout CG Domcke:2020quw .
Within the framework presented in this paper, we demonstrate that the initial conditions required for successful WILG can also be used to address the generation of the DM in our universe Planck:2018vyg ; Cirelli:2024ssz . In particular, if DM has a particle nature and consists primarily of either dark particles or dark anti-particles, then a primordial charge asymmetry in the SM sector can induce suitable initial conditions for ADM. Similarly, a primordial charge asymmetry in the dark sector can also induce initial conditions suitable for WILG. This idea is closely related to the standard ADM paradigm Kaplan:1991ah ; Nussinov:1985xr ; Kaplan:2009ag ; Petraki:2013wwa ; Zurek:2013wia , where the generation of the BAU and the DM abundance are tightly connected. In many ADM models, either a $B-L$ asymmetry or a DM asymmetry is generated at high temperatures, and subsequently transferred to the other sector Shelton:2010ta ; Haba:2010bm ; Buckley:2010ui ; Blennow:2010qp ; Bernal:2016gfn ; Hall:2019ank ; Hall:2021zsk ; Barman:2021tgt ; Asadi:2025vli ; ADMSeesaw . A standard transfer mechanism is provided by a higher-dimensional operator of the form $\mathcal{O}_{B-L}\mathcal{O}_{D}/M^{n}$, where $\mathcal{O}_{B-L}$ violates $B-L$ and $\mathcal{O}_{D}$ violates a global DM charge Zurek:2013wia . At high temperatures, the interaction rates induced by this operator are efficient, and the asymmetry is freely shared between the two sectors. As the temperature in the early universe drops, the interaction decouples, and an asymmetry is separately frozen in the dark and visible sectors. Finally, as the temperature drops below the mass of the DM particle, particle-antiparticle annihilations erase the symmetric component of the DM density, resulting in a DM density determined by the particle-antiparticle asymmetry and prediction of a $\mathcal{O}$(GeV) DM mass Zurek:2013wia .
However, as will be shown in this work, the restrictions on the higher-dimensional interaction transferring asymmetry between the two sectors can be significantly relaxed in the context of the WILG paradigm. Rather than violating $B-L$, it is sufficient if the interaction takes the form
[…]
where $\mathcal{O}_{C}^{\rm{(SM)}}$ violates one of the conserved charges $C$ used in WILG and $\tilde{\Lambda}$ is a heavy mass scale. In general, $\mathcal{O}_{C}^{\rm{(SM)}}$ can be any gauge-invariant operator, but for concreteness, we will consider the case where it is renormalizable, and reduces to a SM Yukawa-type operator. We will also consider the simplest realization possible for the dark operator $\mathcal{O}^{(D)}_{X}=X^{p}$, where $X$ is a complex scalar asymmetric DM candidate, and $\mathcal{O}_{X}^{(D)}$ violates ”$X$-number” charge by $p$ units. Throughout, we assume that $X$ belongs to a dark sector that allows for sufficiently efficient annihilation of the symmetric abundance of $X$ such that the asymmetric component provides the dominant contribution to the DM density, and that the charge-transfer interaction in Eq. (1) provides the dominant source of $X$-number violation.
The framework presented here, which we name Asymgenesis, can be regarded as a flexible and significant extension of the ADM paradigm. In contrast to standard ADM mechanisms, Asymgenesis allows for a separation between the scale of charge-transfer between the dark and the visible sectors and $B-L$ violation. Indeed, Asymgenesis includes the intriguing possibility that the charge-transfer mechanism need not violate $B-L$. Instead, it is sufficient that any of the (global) SM charges that are approximately conserved at high temperatures are violated, allowing for many more options for charge transfer.
The remainder of the paper is organized as follows. We start by illustrating a simple prototype of the asymgenesis framework before we generalize the description to arbitrary initial conditions and temperature regimes. First, we present results that are valid if the charge-transfer interaction induced by the operator in Eq. (1) was in thermal equilibrium at high temperatures. We then go beyond the assumption of thermal equilibrium by deriving and solving Boltzmann equations that describe charge transfer from the SM to the DM sector and vice versa.
Warmup —  We initiate our analysis by examining the transfer of charge between the visible and the dark sector. The charge asymmetry $q_{i}$ of a SM or DM particle species $i$ is defined as, $q_{i}=n_{i}-n_{\bar{i}}=g_{i}\mu_{i}T^{2}/6$, where $n_{i}$ $(n_{\bar{i}})$ denotes the (anti)particle number density of particle $i$ with chemical potential $\mu_{i}$ and multiplicity $g_{i}$. At high temperatures, the interaction described in Eq. (1) may establish thermal equilibrium between the two sectors, relating their chemical potentials $\mu_{i}$ Harvey:1990qw . Further, all SM and DM chemical potentials can be expressed as linear combinations of a set of linearly independent conserved charges $q_{C}=\bar{\mu}_{C}T^{2}/6$, e.g. by employing the linear-algebra formalism developed in the appendix of Ref. Domcke:2020quw . Here, $\bar{\mu}_{C}$ is introduced for notional convenience: $\bar{\mu}_{C}$ and $q_{C}$ are not conjugate to each other in the usual thermodynamical sense, but both describe the conserved charge. In other words, $\bar{\mu}_{C}$ corresponds to $q_{C}$, written in units of a chemical potential. To illustrate the concept of Asymgenesis, we consider a specific example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$, with $L_{e}$, $H$, and $e_{R}$ representing the first-generation lepton doublet, the SM Higgs doublet, and the right-handed electron, respectively. If the higher-dimensional operator is in thermal equilibrium, the following equilibrium condition relating the SM and the dark sector holds,
$p\cdot\mu_{X}=-\mu_{H}-\mu_{e}+\mu_{L_{e}}.$ (2)
While the operator $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ violates right-handed electron number, the interaction Eq. (1) conserves a combination of right-handed electron number and dark number. Consequently, the conservation of the right-handed electron charge, $q_{e}\sim\mu_{e}/T$, at high temperatures Campbell:1992jd ; Cline:1993vv ; Cline:1993bd ; Bodeker:2019ajh is modified as follows
$\frac{\mu_{e}-(2/p)\cdot\mu_{X}}{T}\equiv\frac{\bar{\mu}_{Xe}}{T}=\text{constant},$ (3)
where we used $g_{X}=2$, which we employ throughout this paper. As an example, if the charge-transfer interaction decouples at temperatures of a few 100100100 TeV, then the ADM chemical potential can be written as
$\frac{1422+481p^{2}}{p}\mu_{X}=-711\bar{\mu}_{Xe}-185\bar{\mu}_{\Delta_{e}}+52\left(\bar{\mu}_{\Delta_{\mu}}+\bar{\mu}_{\Delta_{\tau}}\right).$ (4)
The ADM abundance would in this case be determined by two factors: (i) the primordial asymmetries in $q_{e}$ and $q_{X}$, and (ii) the values of the flavored $B-L$ asymmetries, $\Delta_{\alpha}$, at the time of charge-transfer decoupling. If the lightest RHN, $N_{1}$, has a mass around $M_{N}\sim 10^{2-3}$ TeV and decouples while the portal interaction is still efficient, then WILG in the strong wash-in regime would lead to
$\begin{pmatrix}\bar{\mu}_{\Delta_{e}}\\ \bar{\mu}_{\Delta_{\mu}}\\ \bar{\mu}_{\Delta_{\tau}}\end{pmatrix}=\begin{pmatrix}-25\\ 8\\ 8\end{pmatrix}\frac{p^{2}}{\left(66+30p^{2}\right)}\bar{\mu}_{Xe},$ (5)
and both the ADM abundance and the BAU would be fixed entirely by the primordial asymmetry in $q_{Xe}$. The latter could be generated through a CG mechanism that generates nonzero $q_{e}$ and/or nonzero $q_{X}$. At temperatures above $10^{6}$ GeV, additional SM charges are approximately conserved, and can be used as part of Asymgenesis.
Table 1: The numerical coefficients $x_{C}$ (brown) and $y_{C}$ (green), associated with the conserved charges $\bar{\mu}_{C}=q_{C}6/T^{2}$ as defined in Eq. (6) for the scenario with $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ and $\mathcal{O}_{X}^{(D)}=X^{p}$ in different temperature regimes. The first column shows $T_{\rm{dec}}$, which is identified with $T_{B-L}$ $(T_{\rm{eq}})$ for $\bar{\mu}_{B-L}^{\rm eq}$ $\left(\mu_{X}^{\rm eq}\right)$, see text below Eq. (6), and (i)-(v) indicate the following ranges for $T_{\rm{dec}}$ [GeV]: $\quantity(10^{5},10^{6})$, $\quantity(10^{6},10^{9})$, $\quantity(10^{9},10^{11-12})$, $\quantity(10^{11-12},10^{13})$, $\quantity(10^{13},10^{15})$, respectively. The flavor index $\alpha$ in the second column lists the uncorrelated charged-lepton flavors that interact with $N_{1}$ in the relevant temperature range, assuming the SM–DM portal is active (decoupled) in that range. In the latter case, the results for $x_{C}$ reduce to those in standard WILG, c.f. Table II in Domcke:2020quw , which are included in parentheses in column 333 below. The ✗ symbol marks the absence of the corresponding $\bar{\mu}_{C}$ due to an efficient SM interaction. The coefficients $P$ and $P_{e,\tau}$ encode the primordial $q_{e,\mu,\tau}$ asymmetries’ flavor composition relative to the $N_{1}$ washout direction, as detailed in Domcke:2020quw . See text for further details.
$T_{\rm{dec}}$ Index $\alpha$
$\bar{\mu}_{Xe}$ ($\bar{\mu}_{e}$)
$\bar{\mu}_{2B_{1}-B_{2}-B_{3}}$
$\bar{\mu}_{u-d}$
$\bar{\mu}_{d-s}$
$\bar{\mu}_{B_{1}-B_{2}}$
$\bar{\mu}_{\mu}$
$\bar{\mu}_{u-c}$
$\bar{\mu}_{\tau}$
$\bar{\mu}_{d-b}$
$\bar{\mu}_{B}$
$\bar{\mu}_{u}$
$\bar{\mu}_{\Delta_{\perp}}$
$\bar{\mu}_{\Delta_{\alpha}}$
(i) $e,\mu,\tau$ $(e,\mu,\tau)$
$\frac{-3p^{2}}{22+10p^{2}}$ $(\frac{-3}{10})$,$\frac{-711}{1422+481p^{2}}$
[…]
(ii) $e,\mu,\tau$ $(e,\mu,\tau)$
$\frac{-3p^{2}}{38+17p^{2}}$ $\left(\frac{-3}{17}\right)$,$\frac{-1071}{2142+716p^{2}}$
0 (0)(0)( 0 ), 0
$\frac{-16-7p^{2}}{38+17p^{2}}$ $\left(\frac{-7}{17}\right)$, $\frac{45}{2142+716p^{2}}$
[…]
(iii) $e,\mu,\tau$ $(\parallel_{\tau},\tau)$
$\frac{-3p^{2}}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$,$\frac{-1779}{3558+1178p^{2}}$
0 (0)(0)( 0 ),0
$\frac{-6(7+3p^{2})}{62+27p^{2}}$ $\left(\frac{-123}{247}\right)$,$\frac{135}{3558+1178p^{2}}$
$\frac{-4(7+3p^{2})}{62+27p^{2}}$$\left(\frac{-82}{247}\right)$,$\frac{45}{1779+589p^{2}}$
$\frac{21+9p^{2}}{62+27p^{2}}$ $\left(\frac{123}{494}\right)$,$\frac{-135}{7116+2356p^{2}}$
$\frac{-3(2+p^{2})}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$,$\frac{-6}{1779+589p^{2}}$
$\left(\frac{225}{247}\right)$,
[…]
(iv) $e,\parallel_{e}$ ()($\parallel$)( ∥ )
$\frac{-9p^{2}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-38}{76+23p^{2}}$
$\frac{12(3+p^{2})}{102+37p^{2}}$ $\left(\frac{1}{5}\right)$,$\frac{-3}{76+23p^{2}}$
$\frac{-36(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{5}\right)$,$\frac{9}{76+23p^{2}}$
$\frac{-10(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-1}{6}\right)$,$\frac{5}{152+46p^{2}}$
$\frac{-18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{10}\right)$,$\frac{9}{152+46p^{2}}$
$\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-7}{152+46p^{2}}$
$\frac{18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{3}{10}\right)$,$\frac{-9}{152+46p^{2}}$
$\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-7}{152+46p^{2}}$
$\frac{-16(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-4}{15}\right)$,$\frac{4}{76+23p^{2}}$
$\frac{120+46p^{2}}{306+111p^{2}}$ $\left(\frac{23}{90}\right)$,$\frac{23/6}{76+23p^{2}}$
$\frac{60+23p^{2}}{102+37p^{2}}$ $\left(\frac{23}{30}\right)$,$\frac{7}{152+46p^{2}}$
$\frac{[-16,7]}{152+46p^{2}}$
(v) $e,\parallel_{e}$ ()($\parallel$)( ∥ )
$\frac{-p^{2}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$,$\frac{-5}{10+3p^{2}}$
$\frac{4(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$, $\frac{-1}{30+9p^{2}}$
$\frac{-20(3+p^{2})}{42+15p^{2}}$$\left(\frac{-5}{6}\right)$,$\frac{5}{30+9p^{2}}$
$\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$, $\frac{1}{20+6p^{2}}$
$\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$,$\frac{1}{20+6p^{2}}$
$\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$,$\frac{-1}{20+6p^{2}}$
$\frac{2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{1}{4}\right)$,$\frac{-1}{20+6p^{2}}$
$\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$, $\frac{-1}{20+6p^{2}}$
$\frac{-8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{-1}{3}\right)$,$\frac{2}{30+9p^{2}}$
$\frac{10+4p^{2}}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$,$\frac{2}{30+9p^{2}}$
$\frac{8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{3}\right)$, $\frac{-2}{30+9p^{2}}$
$\frac{8+3p^{2}}{14+5p^{2}}$ $\left(\frac{3}{4}\right)$,$\frac{1}{20+6p^{2}}$
[…]
Asymgenesis —  Consider a general scenario where primordial charges in both the visible and the dark sector can take arbitrary values and the charge-transfer interaction is in thermal equilibrium at high temperatures. The $B-L$ asymmetry in the strong wash-in regime and the ADM asymmetry can then, in full generality, be expressed as
$\frac{q_{B-L}^{\rm{win}}}{s}=\sum_{C\neq\Delta_{\alpha}}x_{C}\frac{q_{C}}{s}\Bigg{|}_{T_{B-L}},\quad\frac{q_{X}}{s}=2p\sum_{C}y_{C}\frac{q_{C}}{s}\Bigg{|}_{T_{\text{eq}}},$ (6)
where $T_{B-L}$ $(T_{\rm{eq}})$ denotes the temperature when RHN interactions (the charge-transfer interaction) become(s) inefficient. Here, $\Delta_{\alpha}=B/3-L_{\alpha}$ denote lepton-flavor asymmetries corresponding to the flavor combinations that are decohered at the relevant temperature scale. At temperatures below $T_{\rm{eq}}$, both $q_{e}$ and $q_{X}$ are conserved separately. In particular, for $T_{B-L}<T_{\rm{eq}}$, $C=Xe$ can be replaced with $C=e$ in Eq. (6), with $q_{e}$ given by
$\frac{q_{e}}{s}=\frac{q_{Xe}}{s}+2\sum_{C}y_{C}\frac{q_{C}}{s}\Bigg{|}_{T_{\text{eq}}}.$ (7)
In this case, the numerical values for $x_{C}$ reduce to the result from standard WILG Domcke:2020quw . The values for $x_{C}$ and $y_{C}$ are shown for various temperature regimes in Table 1 for our example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ footnote . The coefficients without parentheses correspond to $T_{\rm{CG}}>T_{B-L}>T_{\rm{eq}}$, while the coefficients in parentheses correspond to $T_{\rm{CG}}>T_{\rm{eq}}>T_{B-L}$, where $T_{\rm{CG}}$ is the temperature scale associated with CG. The table shows that both the asymmetry in $X$ particles and the BAU are typically related to the primordial charge asymmetries generated during CG by an $\mathcal{O}(0.1-1)$ number. The analogous statement in the context of standard ADM models is that if the $B-L$ and $X$-violating higher-dimensional operator is in thermal equilibrium at high temperatures, then the BAU is related to the asymmetric component of DM by a $\mathcal{O}(1)$ number, barring potential washout processes. In conclusion, the framework considered so far predicts DM masses of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, which is similar to standard ADM models.
Beyond thermal equilibrium: SM to DM — Up to this point, we have analyzed the exchange of charges between the SM and the dark sector under the assumption that the charge-transfer interaction attained thermal equilibrium in the early universe. This requirement, however, is not essential. The same dynamics can be treated in an out-of-equilibrium framework using semiclassical Boltzmann equations. We will first consider the scenario where a CG mechanism generates primordial charges in the SM sector, and derive equations describing charge transfer to the DM sector. To facilitate this description, we introduce the following quantities
$Y_{\Delta i}$ $=\frac{q_{i}}{s},\quad Y_{C}=\frac{q_{C}}{s},\quad z=\frac{M_{N}}{T}.$ (8)
We find that the Boltzmann equation governing charge transfer from the SM plasma to the dark sector can be conveniently written as follows,
$\frac{dY_{\Delta X}}{dz}=\left[Y_{\Delta X}^{\rm eq}(z)-Y_{\Delta X}(z)\right]W(z),$ (9)
where
$Y_{\Delta X}^{\rm eq}(z)=2p\sum_{C}y_{C}(z)Y_{C}(z).$ (10)
The washout term, $W(z)$, associated with the interactions induced by Eq. (1) takes the following form,
$W=\frac{\Gamma^{w}}{zH},\quad\Gamma^{w}=\frac{M_{N}^{2n+1}}{z^{2n+1}\Lambda^{2n}},\quad H=\sqrt{\frac{\pi^{2}g_{*}}{90}}\frac{M_{N}^{2}}{z^{2}M_{P}},$ (11)
assuming $X$ is relativistic in the relevant temperature regime. Here, $H$ is the Hubble parameter for a radiation-dominated universe, $g_{*}$ is the number of relativistic degrees of freedom, and $M_{P}$ denotes the reduced Planck mass. Finally, $\Lambda\sim\tilde{\Lambda}$, up to a model-dependent numerical prefactor, where $\tilde{\Lambda}$ is the effective-operator scale introduced in Eq. (1). The solution to Eq. (9) takes the following form
$Y_{\Delta X}(z)$ $=\frac{1}{\omega(z)}\left(Y_{\Delta X}(z_{\text{CG}})+\int_{z_{\text{CG}}}^{z}\omega(\tilde{z})W(\tilde{z})Y_{\Delta X}^{\rm{eq}}(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (12)
where
$\omega(z)$ $=\exp\left(\int_{z_{\text{CG}}}^{z}W(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (13)
and $z_{\rm{CG}}=M_{N}/T_{\rm{CG}}$. As the electron-Yukawa interaction comes into equilibrium, $Y_{\Delta X}^{\rm{eq}}$ goes to zero. If, at that stage, the charge-transfer interaction characterized by $\Gamma^{w}$ remains efficient, Eqs. (12) and (13) show that $Y_{\Delta X}$ is driven to zero, thereby washing out the $X$ asymmetry. Meanwhile, if the charge-transfer interactions are initially in thermal equilibrium and decouple above the electron Yukawa equilibration temperature, $T_{e}\approx 85$ TeV Bodeker:2019ajh , then the final ADM yield, $Y_{\Delta X}^{\rm{end}}$, is driven exponentially close to $Y_{\Delta X}^{\rm{eq}}$, as given by Eq. (10) and Table 1. The attraction of $Y_{\Delta X}^{\rm{end}}$ towards $Y_{\Delta X}^{\rm{eq}}$ driven by $W$, can be referred to as dark wash-in, see also Asadi:2025vli .
Suppose instead that the effective-operator scale is so high that the charge-transfer interaction never enters thermal equilibrium. The charge-transfer then becomes a UV freeze-in (FI) process Giudice:2000ex ; Hall:2009bx ; Elahi:2014fsa . In such a case, $Y_{\Delta X}^{\rm{end}}$ is to a good approximation, determined by the value of $Y_{\Delta X}^{\rm{eq}}$ right after CG, and the value of $\Lambda$ and $T_{\rm{CG}}$. To unify the UV FI regime and the thermal-equilibrium regime, we find it useful to express the final asymmetry in $\Delta X$ as
$Y_{\Delta X}^{\rm{end}}=\kappa(z^{\prime})Y_{\Delta X}^{\rm{eq}}(z^{\prime}),$ (14)
where $z^{\prime}=\text{Max}[z_{\rm{CG}},z_{\text{eq}}]$ and $\kappa$ parametrizes the charge-transfer efficiency. Numerical values for $\kappa$ for the case, $\mathcal{O}_{X}^{(D)}=X^{2}$, are shown in Fig. 1, where the horizontal and vertical axes show $T_{\rm{CG}}$ and $\Lambda$, respectively, and the black solid curve shows $H(z_{\rm{CG}})=\Gamma^{w}(z_{\rm{CG}})$. The latter marks the transition between the two regimes: The area below the curve corresponds to the thermal-equilibrium regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})>1$, where $\kappa\simeq 1$ and the asymmetry is well described by Tab. 1. The area above the curve shows the UV FI regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})<1$, and in the limit $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})\ll 1$, we have $\kappa\simeq\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})$. For charge-transfer from the SM to the DM sector, the asymmetry in the dark sector is significantly smaller than the BAU in large portions of parameter space where Asymgenesis proceeds via UV FI. The black-dashed curve shows the contour corresponding to $\kappa=10^{-3}$, which would point to DM masses of $\mathcal{O}(1)$ TeV. This roughly represents an upper bound on ADM masses when the symmetric annihilation cross section is required to be large enough for the ADM component to be the dominant contribution to the DM relic density while respecting perturbative unitarity Griest:1989wd ; Cohen:2010kn ; Petraki:2013wwa . Interestingly, for $T_{\rm{CG}}=10^{14-15}$ GeV, it could be possible to account for the DM abundance even if the scale of NP associated with charge transfer is related to the scale of grand unification or possibly quantum-gravity effects, c.f. the upper-right corner of the figure. In summary, the region below the black solid curve in Fig. 1 corresponds to DM masses in the range of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, while the area below the black dashed curve corresponds to DM masses up to $\mathcal{O}(1)$ TeV. Above the black dashed curve, the symmetric component cannot annihilate efficiently due to the constraints imposed by perturbative unitarity.
Figure 1: Numerical result for $\kappa$ $(\tilde{\kappa})$, defined in Eq. (14) (Eq. (19)) for charge transfer from the SM (DM) sector to the DM (SM) sector, in the $T_{\rm{CG}}-\Lambda$ plane. See text for details.
Beyond thermal equilibrium: DM to SM —  Next, we consider the final possibility in which the initial conditions are set by a CG mechanism in the dark sector, and the effective operator scale is sufficiently high that the charge-transfer interaction never reaches thermal equilibrium. For the example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$, we obtain a Boltzmann equation for the charge transfer to the SM sector of the following form,
$\frac{dY_{\Delta e}}{dz}=\left[Y_{\Delta e}^{\rm{eq}}(z)-Y_{\Delta e}(z)\right]\widetilde{W}(z),$ (15)
where
$Y_{\Delta e}^{\rm{eq}}(z)=Y_{Xe}(z)+2\sum_{C}y_{C}Y_{C}(z),$ (16)
and $\widetilde{W}(z)$ takes the same form as in Eq. (11) although the numerical value of the scale $\Lambda$ entering $\widetilde{W}$ might differ from that entering $W$ by an $\mathcal{O}(1)$ number. In an abuse of notation, we will denote both the heavy mass scale entering $W$ and $\widetilde{W}$ by $\Lambda$ in our results. With this, the solution to Eq. (15) reads
$Y_{\Delta e}(z)$ $=\frac{1}{\tilde{\omega}(z)}\left(Y_{\Delta e}(z_{\text{CG}})+\int_{z_{\text{CG}}}^{z}\tilde{\omega}(\tilde{z})\widetilde{W}(\tilde{z})Y_{\Delta e}^{\rm{eq}}(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (17)
where
$\tilde{\omega}(z)$ $=\exp\left(\int_{z_{\text{CG}}}^{z}\widetilde{W}(\tilde{z})\,\mathrm{d}\tilde{z}\right).$ (18)
We can express the resulting right-handed electron asymmetry as
$Y_{\Delta e}^{\rm{end}}=\tilde{\kappa}(z^{\prime})Y_{\Delta e}^{\rm{eq}}(z^{\prime}),$ (19)
in full analogy with Eq. (14). Written in this form, with our abuse of notation for $\Lambda$, the results for $\tilde{\kappa}$ are equal to $\kappa$, which are displayed in Fig. 1. The parameter $\tilde{\kappa}$ indirectly parametrizes the outcome of WILG in the strong wash-in regime, which takes the following form in the scenario considered in this section
$\frac{q_{B-L}^{\rm{win}}}{s}=x_{e}(z_{B-L})\tilde{\kappa}(z^{\prime})Y_{\Delta e}^{\rm{eq}}(z^{\prime}).$ (20)
In the scenario where the portal interaction attains thermal equilibrium, we have $Y_{\Delta e}^{\rm{eq}}(z^{\prime})\sim Y_{Xe}\sim Y_{\Delta X}(z^{\prime})$. This condition implies that, if the BAU is explained via WILG, the ADM mass must lie in the range $\mathcal{O}(1)-\mathcal{O}(10)$ GeV. In the FI regime, where $\tilde{\kappa}\ll 1$, the final ADM yield is typically much larger than the BAU, and one would expect ADM masses in the range $\mathcal{O}(\tilde{\kappa})-\mathcal{O}(10\,\tilde{\kappa})$ GeV. Specifically, the black dashed line in Fig. 1 corresponds to DM masses in the MeV range.
Finally, we note that since Asymgenesis builds on the type-I seesaw extension of the SM, it might be interesting to apply the framework to BSM models with gauged $U(1)_{B-L}$ symmetry. Contrary to conventional ADM models, Asymgenesis can accommodate scenarios where the charge-transfer mechanism decouples at temperatures that are orders of magnitude above the $B-L$-breaking scale, and still result in comparable abundances of baryonic matter and ADM.
Conclusion — In this work, we have proposed Asymgenesis, a framework based on the type-I seesaw model that can be used to account for the BAU and the observed DM density from initial conditions appropriate to either ADM or WILG. Compared to standard ADM scenarios that rely on higher-dimensional operators, Asymgenesis offers significantly greater flexibility in the charge-transfer interactions between the dark and visible sectors. Crucially, it disentangles the temperature at which the charge-transfer interactions decouple from the temperature scale at which $B-L$ violation becomes effective. By combining the advantages of WILG and ADM, the Asymgenesis framework considerably widens the model-building landscape for asymmetric dark matter.
Acknowledgments — We are grateful to Kai Schmitz for insightful discussions and for pointing out an incorrect factor in Eq. (3) in the first arXiv version. We thank Mathias Becker and Nicholas Leister for comments on the draft. M. A. M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Centre “Neutrinos and Dark Matter in Astro- and Particle Physics” (SFB 1258). The work of M. A. M. and S. W. was supported by the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the German Excellence Strategy (Project No. 390831469).